find the distance between the points (4,4) and (-7- -7) round to the nearest tenth
Answer by swincher4391(1107)

sqrt((x2-x1)^2 + (y2-y1)^2) = sqrt((4+7)^2 + (4+7)^2) = srqt(242) = 15.6